As sales manager, Joe Batista was given the following static budget report for selling expenses in the Clothing Department of Soria Company for the month of October. SORIA COMPANY Clothing Department Budget Report For the Month Ended October 31, 2020 Difference Budget Actual Favorable Unfavorable Neither Favorable nor Unfavorable Sales in units 7,900 9,000 1,100 Favorable Variable expenses Sales commissions $1,896 $2,430 $534 Unfavorable Advertising expense 1,106 630 476 Favorable Travel expense 3,318 3,150 168 Favorable Free samples given out 1,185 990 195 Favorable Total variable 7,505 7,200 305 Favorable Fixed expenses Rent 1,200 1,200 –0– Neither Favorable nor Unfavorable Sales salaries 1,000 1,000 –0– Neither Favorable nor Unfavorable Office salaries 700 700 –0– Neither Favorable nor Unfavorable Depreciation—autos (sales staff) 400 400 –0– Neither Favorable nor Unfavorable Total fixed 3,300 3,300 –0– Neither Favorable nor Unfavorable Total expenses $10,805 $10,500 $305 Favorable As a result of this budget report, Joe was called into the president’s office and congratulated on his fine sales performance. He was reprimanded, however, for allowing his costs to get out of control. Joe knew something was wrong with the performance report that he had been given. However, he was not sure what to do, and comes to you for advice. Prepare a budget report based on flexible budget data to help Joe. (List variable costs before fixed costs.) SORIA COMPANY Selling Expense Flexible Budget Report Clothing Department For the Month Ended October 31, 2020 Difference Budget Actual Favorable Unfavorable Neither Favorable nor Unfavorable $ $ $ $ $ $.
Answer #1
SORIA COMPANY
Selling Expense Flexible Budget Report
Clothing Department
For the Month Ended October 31, 2020
Difference
Budget

Actual

Favorable

Unfavorable

Neither Favorable Nor Unfavorable

Sales in units 9000 9000
Variable expenses::
Sales commissions (1896/7900*9000)= 2160 $2430 (2160-2430)= 270 Unfavorable
Advertising expense (1106/7900*9000)= 1260 630 (1260-630)= 630 Favorable
Travel expense (3318/7900*9000)= 3780 3150 (3780-3150= 630 Favorable
Free samples (1185*7900/9000)= 1350 990 (1350-990)= 360 Favorable
Total variable expenses 8550 7200 (8550-7200)= 1350 Favorable
Fixed expenses:
Rent 1200 1200 0 NONE
Sales salaries 1000 1000 0 NONE
Office salaries 700 700 0 NONE
Depreciation—autos (sales staff) 400 400 0
Total fixed expenses 3300 3300 0 NONE
Total expenses $11850 $10500 $1350 Favorable

Joe should not been reprimanded as the variable cost of the actual budget is $1350 lower than the variable cost of the budget report. So, Joe should not reprimanded.